OpenStudy (anonymous):

OK so how do I plug this into the calculator? Im using a TI-83 Plus. f(x) = sin x - 5

Parth (parthkohli):

What exactly do you have to do here?

OpenStudy (anonymous):

well the question just wants me to graph it.

OpenStudy (amistre64):

hit the y= button onthe top left

Parth (parthkohli):

If there is a question like "find \(f(x)\) when \(x = 3\) where \(f(x) = \sin x - 5\)," then you may enter:\[\sin (3) - 5 \]

OpenStudy (anonymous):

nooo its more like Graphing Trigonometric Functions

OpenStudy (amistre64):

do you see the y= button?

OpenStudy (amistre64):



OpenStudy (anonymous):

yea i see the y= button and stuff i just don't know exactly what to right. okok heres the problem...

OpenStudy (amistre64):

hit the buttons: Y= SIN (X,T,O,n) - 5 GRAPH

OpenStudy (anonymous):

ohhhhhh thanks! ha i feel so dumb lol

OpenStudy (amistre64):

the (X,T,O,n) button is the variable button ... i dont know why they decided to name it that other than to confuse me ;)

OpenStudy (amistre64):

if the first y= slot already has stuff in it, you would want to hit CLEAR to clean it out before entering the sinx-5 stuff

OpenStudy (anonymous):

and i also just realized my calculator was in the wrong mode! thanks again ha.

OpenStudy (amistre64):

your welcome; oh, and i just tried it on mine, the sin function wants you to end it with the ")" before tacking on the -5 part

OpenStudy (anonymous):

Great picture of a calculator there, amistre!

OpenStudy (amistre64):

i was wanting to draw the mr professor one i had as a kid :)